OpenStudy (anonymous):

if your given the radius 6 meters and area of sector of circle 8 squared meters how do you find the central angle?

OpenStudy (anonymous):

sector area is a/360 * pi r^2

OpenStudy (anonymous):

the formula is central angle/2pi = A /pir^2.

OpenStudy (anonymous):

so what he said above me.

OpenStudy (anonymous):

you're *

OpenStudy (anonymous):

8 = a/360 36pi 113a = 8 a = 14.125/360 so is the measure 14.125 degrees. i've no idea. lol this is weird.

OpenStudy (anonymous):

a/2pi = 8/36pi a/pi = 4/36pi 12.56 = 113a a = 9 according to marcs formula so yeah :)

OpenStudy (anonymous):

say thanks to my cal teacher haha just came from the unit circle :) but your definition looked right you were just mising the area of the circle i think?

OpenStudy (anonymous):

Yep, well i'm only in geometry and we did stuff like this. but it's different working backwards.

OpenStudy (anonymous):

Id say more here but this isnt a conversation area :) haha